OpenStudy (anonymous):

Mikalee, a fighter pilot, ejects from her burning plane. Once she is safely on the ground, she launches a flare to attract the attention of a rescue plane. The initial velocity of the flare is 1600 ft/sec. The path of the flare is represented by the function h(t) = - 16t^2 + 1600t where t is the time in seconds and h(t) is the height of the flare. How long does it take the flare to reach the maximum height? 40 sec 4 sec 50 sec 40,000 sec

OpenStudy (anonymous):

\[ h(t) = - 16t^2 + 1600t\] max is at \(t=-\frac{b}{2a}=-\frac{1600}{2\times 16}\)